OpenStudy (anonymous):

9 over 4- 8 over 12=?

OpenStudy (whpalmer4):

is that \[\frac{9}{4} - \frac{8}{12}\]?

OpenStudy (chris911):

reduce first

OpenStudy (chris911):

right @whpalmer4 ??

OpenStudy (anonymous):

yes to palmer

OpenStudy (whpalmer4):

No, can't reduce...

OpenStudy (whpalmer4):

can't reduce in a useful fashion, that is.

OpenStudy (whpalmer4):

need to make a common denominator. 12 is a multiple of 4, so we can use 12 as the common denominator. What do we need to do to make the denominator of \(\dfrac{9}{4}\) be \(12\)?

OpenStudy (anonymous):

I know there is 27-8 over 12

OpenStudy (whpalmer4):

Yes, \[\frac{9}{4}*\frac{3}{3}-\frac{8}{12} = \frac{27}{12} - \frac{8}{12} = \frac{27-8}{12} = \]

OpenStudy (anonymous):

19 over 12

OpenStudy (whpalmer4):

that's correct!

OpenStudy (anonymous):

Is that the answer tho?

OpenStudy (whpalmer4):

Yes, unless the problem asks that it be made a proper fraction. You can't reduce 19/12, as 19 is a prime number, and so has no factors other than 1 and 19.

OpenStudy (anonymous):

 actually this is the correct meaning of the question...

OpenStudy (anonymous):

It was 19/12

OpenStudy (whpalmer4):

alternative answer would be \[1\frac{7}{12}\]

OpenStudy (anonymous):

no answer would be -27

OpenStudy (anonymous):

see this you will get it...

OpenStudy (whpalmer4):

@chetan552 she knows what the problem looks like on her screen or in the book. yes, what you provided was the correct interpretation of what she typed, but she didn't type what she meant.

OpenStudy (whpalmer4):

She even told us that the correct answer was 19/12...

OpenStudy (whpalmer4):

Part of the joy of OpenStudy is that you have to decipher what the problem really is from what may be an imperfectly transmitted description :-)